# “Whose Side Are You On?” Estimating Ideology of Political and News Content Using Large Language Models and Few-shot Demonstration Selection

### v2

@@ 6 Conclusion
We explored the efficacy of few-shot learning using LLMs for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance in the few-shot setting, surpassing traditional supervised baselines. By leveraging set-cover based ICL, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
@@ 7 Limitations
Despite the promising results and contributions outlined in this paper, several limitations must be acknowledged to identify areas for future work. The performance of our approach is heavily reliant on the capabilities of pre-trained LLMs such as GPT-4o, Llama2-13B, and Mistral-7B. Variability in model updates or differences in underlying pre-training data could affect reproducibility and consistency across different deployments. In addition, while incorporating metadata, such as source information improved performance, it also introduced the potential for biased predictions. Models may disproportionately rely on known ideological leanings of sources, potentially misclassifying nuanced or neutral content.
Furthermore, our experiments focus on text, and extending these methods to classify ideology in non-textual data such as video or audio remains an open challenge and an important direction for future work. Lastly, we acknowledge that the datasets used for evaluation, while diverse, may not fully capture the breadth of content and ideological nuance present in real-world social media platforms. For instance, the balance of liberal, neutral, and conservative content in the datasets may not reflect natural distributions, and also the majority of content consumed by users is generated by other users themselves, not by news media organizations or YouTube channels, suggesting the need to adapt our approach to user-generated posts and comments.
@@ 8 Ethics Statement
Automated ideology estimation carries inherent risks, including misuse for surveillance or censorship. Furthermore, reliance on LLMs can perpetuate biases present in training data, which might influence predictions in unintended ways. While the study evaluates multiple datasets, the generalizability of the approach across all social media platforms—with varying content types and user behaviors–—is yet to be fully established.
@@
Appendix A Prompts used for LLMs
@@
Appendix B Ablations on different parameters
B.1 MLP Baseline
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.51 | 0.38 | 0.43 | ||
| ✓ | 0.61 | 0.55 | 0.69 | ||
| ✓ | 0.55 | 0.38 | 0.47 | ||
| ✓ | ✓ | 0.63 | 0.62 | 0.57 | |
| ✓ | ✓ | 0.57 | 0.43 | 0.46 | |
| ✓ | ✓ | 0.62 | 0.60 | 0.58 | |
| ✓ | ✓ | ✓ | 0.64 | 0.61 | 0.52 |
B.2 RoBERTa baseline
| YouTube | Ad Fontes | AllSides | |||
| Title | Source | Desc. | Slant | News | News |
| ✓ | 0.57 | 0.35 | 0.52 | ||
| ✓ | 0.45 | 0.48 | 0.48 | ||
| ✓ | 0.62 | 0.35 | 0.50 | ||
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | 0.64 | 0.38 | 0.53 | |
| ✓ | ✓ | 0.65 | 0.44 | 0.58 | |
| ✓ | ✓ | ✓ | 0.65 | 0.44 | 0.55 |
B.3 GPT-4o
Table 3 shows the accuracy of the GPT-4o model under different ablations of title, source, and description across the three datasets and varying number of demonstrations considered. We see that the model performs the best when given a higher number of demonstrations for the YouTube slant and Ad Fontes news datasets. For the YouTube dataset, the model performed best when title, source, and description were all provided whereas for the Ad Fontes news dataset, only title and source were sufficient. The AllSides news dataset showed a different trend in that zero demonstrations gave the best score when only the source was provided. We observed in the test set for this dataset that source and target label were highly correlated (i.e., a video from Fox, a well-known conservative outlet, was almost always labeled conservative) making it simpler for the model to just assign the source ideology as the target ideology. In Appendix E, we investigate this further by considering edge cases in the test set.
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.64 | 0.69 | 0.50 | ||
| 0 | ✓ | 0.63 | 0.80 | 0.88 | ||
| 0 | ✓ | 0.56 | 0.64 | 0.44 | ||
| 0 | ✓ | ✓ | 0.64 | 0.78 | 0.74 | |
| 0 | ✓ | ✓ | 0.63 | 0.75 | 0.55 | |
| 0 | ✓ | ✓ | 0.59 | 0.82 | 0.85 | |
| 0 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.72 |
| 4 | ✓ | 0.67 | 0.74 | 0.55 | ||
| 4 | ✓ | ✓ | 0.73 | 0.87 | 0.83 | |
| 4 | ✓ | ✓ | 0.69 | 0.80 | 0.58 | |
| 4 | ✓ | ✓ | ✓ | 0.72 | 0.86 | 0.82 |
| 8 | ✓ | 0.68 | 0.73 | 0.55 | ||
| 8 | ✓ | ✓ | 0.72 | 0.87 | 0.86 | |
| 8 | ✓ | ✓ | 0.70 | 0.81 | 0.59 | |
| 8 | ✓ | ✓ | ✓ | 0.71 | 0.86 | 0.84 |
| 12 | ✓ | 0.69 | 0.74 | 0.55 | ||
| 12 | ✓ | ✓ | 0.74 | 0.87 | 0.87 | |
| 12 | ✓ | ✓ | 0.71 | 0.82 | 0.59 | |
| 12 | ✓ | ✓ | ✓ | 0.71 | 0.87 | 0.84 |
B.4 Llama2
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.53 | 0.49 | 0.42 | ||
| 0 | ✓ | 0.61 | 0.52 | 0.82 | ||
| 0 | ✓ | 0.52 | 0.49 | 0.37 | ||
| 0 | ✓ | ✓ | 0.63 | 0.57 | 0.65 | |
| 0 | ✓ | ✓ | 0.58 | 0.57 | 0.46 | |
| 0 | ✓ | ✓ | 0.48 | 0.65 | 0.71 | |
| 0 | ✓ | ✓ | ✓ | 0.58 | 0.69 | 0.59 |
| 4 | ✓ | 0.64 | 0.60 | 0.54 | ||
| 4 | ✓ | ✓ | 0.68 | 0.79 | 0.75 | |
| 4 | ✓ | ✓ | 0.66 | 0.68 | 0.51 | |
| 4 | ✓ | ✓ | ✓ | 0.64 | 0.79 | 0.68 |
| 8 | ✓ | 0.60 | 0.56 | 0.51 | ||
| 8 | ✓ | ✓ | 0.63 | 0.76 | 0.74 | |
| 8 | ✓ | ✓ | 0.63 | 0.67 | 0.52 | |
| 8 | ✓ | ✓ | ✓ | 0.60 | 0.78 | 0.67 |
| 12 | ✓ | 0.60 | 0.60 | 0.50 | ||
| 12 | ✓ | ✓ | 0.62 | 0.75 | 0.72 | |
| 12 | ✓ | ✓ | 0.58 | 0.68 | 0.49 | |
| 12 | ✓ | ✓ | ✓ | 0.56 | 0.78 | 0.67 |
B.5 Mistral
| YouTube | Ad Fontes | AllSides | ||||
| Title | Source | Desc. | Slant | News | News | |
| 0 | ✓ | 0.61 | 0.57 | 0.44 | ||
| 0 | ✓ | 0.61 | 0.72 | 0.83 | ||
| 0 | ✓ | 0.52 | 0.59 | 0.34 | ||
| 0 | ✓ | ✓ | 0.67 | 0.72 | 0.71 | |
| 0 | ✓ | ✓ | 0.62 | 0.62 | 0.41 | |
| 0 | ✓ | ✓ | 0.57 | 0.73 | 0.66 | |
| 0 | ✓ | ✓ | ✓ | 0.65 | 0.74 | 0.58 |
| 4 | ✓ | 0.61 | 0.57 | 0.42 | ||
| 4 | ✓ | ✓ | 0.65 | 0.73 | 0.73 | |
| 4 | ✓ | ✓ | 0.62 | 0.61 | 0.44 | |
| 4 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.60 |
| 8 | ✓ | 0.62 | 0.59 | 0.44 | ||
| 8 | ✓ | ✓ | 0.64 | 0.71 | 0.73 | |
| 8 | ✓ | ✓ | 0.63 | 0.62 | 0.48 | |
| 8 | ✓ | ✓ | ✓ | 0.63 | 0.68 | 0.61 |
| 12 | ✓ | 0.64 | 0.61 | 0.44 | ||
| 12 | ✓ | ✓ | 0.66 | 0.71 | 0.73 | |
| 12 | ✓ | ✓ | 0.61 | 0.63 | 0.48 | |
| 12 | ✓ | ✓ | ✓ | 0.63 | 0.69 | 0.61 |
Appendix C Candidate set size can be smaller
| LLM | YouTube Slant | Ad Fontes News | AllSides News |
| GPT-4o | 0.66 | 0.74 | 0.52 |
| Llama2 | 0.60 | 0.57 | 0.50 |
| Mistral | 0.63 | 0.59 | 0.42 |
Appendix D Balanced demonstration selection is better than random
| Dataset | Random | Balanced BERTScore |
| YouTube Slant | 0.52 | 0.61 |
| Ad Fontes News | 0.49 | 0.57 |
| AllSides News | 0.39 | 0.42 |
Appendix E Sources can be misleading
Our prior results showed that providing the source information significantly increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart222https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.

### v3

@@ 5 Conclusion
This study explored the efficacy of few-shot learning using Large Language Models for the task of ideology estimation across diverse textual datasets, including news articles and YouTube videos. Our findings demonstrate that LLMs, particularly GPT-4o, achieve state-of-the-art performance, surpassing traditional supervised baselines. By leveraging advanced few-shot in-context learning methods, we addressed challenges of scalability and robustness, showcasing the ability of LLMs to adapt to evolving contexts with minimal supervision.
@@ 6 Limitations
Despite the promising results and contributions outlined in this paper, several limitations must be acknowledged to identify areas for future work. The performance of our approach is heavily reliant on the capabilities of pre-trained LLMs such as GPT-4o, Llama2-13B, and Mistral-7B. Variability in model updates or differences in underlying pre-training data could affect reproducibility and consistency across different deployments. In addition, while incorporating metadata, such as source information improved performance, it also introduced the potential for biased predictions. Models may disproportionately rely on known ideological leanings of sources, potentially misclassifying nuanced or neutral content. This issue might be further exacerbated by the pre-processing guardrails that serve to filter data, but inadvertently possess their own robustness and bias concerns (Achara and Chhabra, 2025; Yang et al., 2024).
Furthermore, our experiments focus on text, and extending these methods to classify ideology in non-textual data such as video or audio remains an open challenge and an important direction for future work. Lastly, we acknowledge that the datasets used for evaluation, while diverse, may not fully capture the breadth of content and ideological nuance present in real-world social media platforms. For instance, the balance of liberal, neutral, and conservative content in the datasets may not reflect natural distributions, and also the majority of content consumed by users is generated by other users themselves, not by news media organizations or YouTube channels, suggesting the need to adapt our approach to user-generated posts and comments. Finally, while beyond the scope of our work, LLMs can themselves have several robustness issues (Askari et al., 2025; Chhabra et al., 2024) that can result in potential downstream evaluation issues.
@@ 7 Ethics Statement
Automated ideology estimation carries inherent risks, including misuse for surveillance or censorship. Furthermore, reliance on LLMs can perpetuate biases present in training data, which might influence predictions in unintended ways. While the study evaluates multiple datasets, the generalizability of the approach across all social media platforms, with varying content types and user behaviors, is yet to be fully established.
@@
Appendix A Prompts used for the LLM
@@
Appendix B Statistical significances
We use the McNemar test to determine the statistical differences between the accuracies of two specific classifiers. Tables 4, 5, and 6 show the significance values for the models trained on the YouTube Slant, Ad Fontes news, and AllSides news datasets respectively.
| GPT | Llama | Mistral | |||||||||||
| T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | ||
| GPT | T | — | 0.13 | 0.06 | 33.06** | 34.18** | 0.21 | 11.05** | 10.75** | 4.50* | 4.21* | 2.34 | 0.39 |
| T/S | 0.13 | — | 0.52 | 27.95** | 35.79** | 0.69 | 12.10** | 12.30** | 4.72* | 2.79 | 3.31 | 0.04 | |
| T/B | 0.06 | 0.52 | — | 42.29** | 30.30** | 0.06 | 10.25** | 10.16** | 2.96 | 5.33* | 1.70 | 0.88 | |
| T/S/B | 33.06** | 27.95** | 42.29** | — | 89.28** | 22.45** | 59.74** | 58.72** | 41.44** | 7.87** | 45.75** | 22.90** | |
| Llama | T | 34.18** | 35.79** | 30.30** | 89.28** | — | 37.09** | 9.03** | 8.62** | 19.19** | 58.08** | 23.16** | 43.91** |
| T/S | 0.21 | 0.69 | 0.06 | 22.45** | 37.09** | — | 7.81** | 8.57** | 1.75 | 5.46* | 0.56 | 1.20 | |
| T/B | 11.05** | 12.10** | 10.25** | 59.74** | 9.03** | 7.81** | — | 0.00 | 1.71 | 26.06** | 4.14* | 16.75** | |
| T/S/B | 10.75** | 12.30** | 10.16** | 58.72** | 8.62** | 8.57** | 0.00 | — | 1.83 | 26.38** | 4.38* | 17.32** | |
| Mistral | T | 4.50* | 4.72* | 2.96 | 41.44** | 19.19** | 1.75 | 1.71 | 1.83 | — | 18.21** | 0.43 | 7.90** |
| T/S | 4.21* | 2.79 | 5.33* | 7.87** | 58.08** | 5.46* | 26.06** | 26.38** | 18.21** | — | 13.56** | 2.79 | |
| T/B | 2.34 | 3.31 | 1.70 | 45.75** | 23.16** | 0.56 | 4.14* | 4.38* | 0.43 | 13.56** | — | 9.75** | |
| T/S/B | 0.39 | 0.04 | 0.88 | 22.90** | 43.91** | 1.20 | 16.75** | 17.32** | 7.90** | 2.79 | 9.75** | — | |
| GPT | Llama | Mistral | |||||||||||
| T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | ||
| GPT | T | — | 34.40** | 20.11** | 449.51** | 96.77** | 32.68** | 38.99** | 224.78** | 51.50** | 2.29 | 18.13** | 288.68** |
| T/S | 34.40** | — | 6.67** | 435.86** | 200.31** | 132.25** | 140.91** | 218.99** | 136.48** | 19.91** | 82.16** | 265.40** | |
| T/B | 20.11** | 6.67** | — | 440.16** | 163.22** | 81.73** | 103.88** | 233.55** | 104.43** | 3.12 | 62.26** | 283.78** | |
| T/S/B | 259.15** | 373.77** | 318.02** | — | 68.03** | 125.03** | 124.95** | 107.43** | 130.00** | 286.31** | 175.39** | 83.12** | |
| Llama | T | 96.77** | 200.31** | 163.22** | 433.77** | — | 21.99** | 21.16** | 236.90** | 18.25** | 138.91** | 40.15** | 276.51** |
| T/S | 32.68** | 132.25** | 81.73** | 475.86** | 21.99** | — | 0.00 | 274.00** | 0.00 | 85.23** | 5.80* | 299.13** | |
| T/B | 38.99** | 140.91** | 103.88** | 468.49** | 21.16** | 0.00 | — | 271.24** | 0.00 | 81.82** | 8.23** | 299.39** | |
| T/S/B | 252.24** | 383.38** | 327.86** | 107.43** | 79.93** | 148.70** | 144.11** | — | 136.57** | 313.88** | 184.20** | 7.11** | |
| Mistral | T | 51.50** | 136.48** | 104.43** | 451.99** | 18.25** | 0.00 | 0.00 | 237.93** | — | 86.05** | 9.88** | 282.52** |
| T/S | 2.29 | 19.91** | 3.12 | 465.55** | 138.91** | 85.23** | 81.82** | 249.35** | 86.05** | — | 33.00** | 295.92** | |
| T/B | 18.13** | 82.16** | 62.26** | 453.61** | 40.15** | 5.80* | 8.23** | 240.67** | 9.88** | 33.00** | — | 295.82** | |
| T/S/B | 240.52** | 343.16** | 295.33** | 83.12** | 56.78** | 104.94** | 107.75** | 7.11** | 113.84** | 263.56** | 161.16** | — | |
| GPT | Llama | Mistral | |||||||||||
| T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | T | T/S | T/B | T/S/B | ||
| GPT | T | — | 277.85** | 32.27** | 255.80** | 23.89** | 60.47** | 6.33* | 25.05** | 17.31** | 186.27** | 47.23** | 33.34** |
| T/S | 277.85** | — | 195.62** | 3.90* | 258.18** | 26.08** | 212.25** | 81.25** | 272.09** | 4.23* | 357.08** | 139.18** | |
| T/B | 32.27** | 195.62** | — | 192.83** | 56.81** | 26.50** | 30.13** | 3.88* | 52.94** | 118.53** | 107.43** | 4.20* | |
| T/S/B | 255.80** | 3.90* | 192.83** | — | 240.60** | 17.80** | 200.62** | 70.24** | 253.02** | 0.56 | 342.79** | 124.91** | |
| Llama | T | 23.89** | 258.18** | 56.81** | 240.60** | — | 247.06** | 12.31** | 132.97** | 2.17 | 260.91** | 0.13 | 80.77** |
| T/S | 60.47** | 26.08** | 26.50** | 17.80** | 247.06** | — | 164.01** | 24.63** | 162.56** | 17.64** | 151.72** | 16.56** | |
| T/B | 6.33* | 212.25** | 30.13** | 200.62** | 12.31** | 164.01** | — | 140.34** | 2.27 | 204.32** | 10.70** | 50.39** | |
| T/S/B | 25.05** | 81.25** | 3.88* | 70.24** | 132.97** | 24.63** | 140.34** | — | 89.82** | 71.34** | 108.30** | 0.17 | |
| Mistral | T | 17.31** | 272.09** | 52.94** | 253.02** | 2.17 | 162.56** | 2.27 | 89.82** | — | 292.91** | 4.31* | 87.31** |
| T/S | 186.27** | 4.23* | 118.53** | 0.56 | 260.91** | 17.64** | 204.32** | 71.34** | 292.91** | — | 339.67** | 138.56** | |
| T/B | 47.23** | 357.08** | 107.43** | 342.79** | 0.13 | 151.72** | 10.70** | 108.30** | 4.31* | 339.67** | — | 175.10** | |
| T/S/B | 33.34** | 139.18** | 4.20* | 124.91** | 80.77** | 16.56** | 50.39** | 0.17 | 87.31** | 138.56** | 175.10** | — | |
Appendix C Sources can be misleading
Our prior results showed that providing the source information substantially increases the accuracy. However, not all content from a given source may share the ideology of the source itself Haroon et al. (2023), as some liberal articles can come from conservative outlets. To highlight this particular scenario, we identify cases in our test set where the ideology of the article or video is different from the ideology of its source for all three datasets (e.g., a liberal article from Fox News or a conservative article from MSNBC). We rely on AllSides’ media bias chart666https://www.allsides.com/media-bias/media-bias-chart for source ideology and map these sources to their corresponding news articles and YouTube videos. We then compare the accuracy of GPT-4o when 1) only title is provided, 2) only source is provided, and 3) both title and source are provided.